OpenStudy (anonymous):

y=sq.rt.(x-4). Find domain and range.

OpenStudy (anonymous):

\[y=\sqrt{x-4}\]

OpenStudy (lgbasallote):

hint for domain: the radicand has to be greater than or equal to zero so do \[x - 4 \ge 0\] then solve for x

OpenStudy (lgbasallote):

does that help?

OpenStudy (anonymous):

so the domain is just \[x \ge 4\]?

OpenStudy (anonymous):

Yes that means that x can be anything equal to or greater than 4. The reason for this is because we can't have a negative value inside the square root. This will give us an imaginary number

OpenStudy (anonymous):

The lowest number that x can be is 4 right? What do you get for y when you plug in x=4 into\[y=\sqrt{x−4}\]

OpenStudy (anonymous):

0

OpenStudy (lgbasallote):

so anyway...that was the domain...

OpenStudy (lgbasallote):

now to find the range...solve for x

OpenStudy (lgbasallote):

do you know how to? or do you need help?

OpenStudy (anonymous):

great we know that the lowest value of y is 0 and we know that we can plug any value greater than 4 until infinity so the range is from \[0 \rightarrow \infty \]

OpenStudy (anonymous):

This means y can be any number as long as it's positive. I don't know the proper notation for this can you help out @lgbasallote

OpenStudy (lgbasallote):

hmm for a proper solution to find the range...solve for x start by squaring both sides \[y^2 = x - 4\] add 4 to both sides \[y^2 + 4 = x\] since there are no restrictions for this term, the range is actually from \(-\infty\) to \(\infty\) <--that means all real values

OpenStudy (lgbasallote):

since there are no restrictions for this equation*

OpenStudy (anonymous):

I thought it was restricted.....from the domain

OpenStudy (lgbasallote):

well square roots have positive and negative roots right?

OpenStudy (anonymous):

the domain says that values of x have to be greater than 4

OpenStudy (anonymous):

no you can't have a negative outcome from a square root.

OpenStudy (lgbasallote):

that's why it's wiser to *solve* for x to avoid confusion hehe

OpenStudy (lgbasallote):

\[\sqrt 4 = \pm 2\]

OpenStudy (lgbasallote):

because 2^2 is 4 and (-2)^2 is also 4

OpenStudy (lgbasallote):

the positive root is just the prinicipal root

OpenStudy (anonymous):

when you graph it, the line starts at (4,0) and continues onward (infinite). so wouldn't it be \[(0, \infty)\]?

OpenStudy (anonymous):

what number do you plug in for x in order to get -2?

OpenStudy (lgbasallote):

if you plug in 8 you will get \[\sqrt 4\] that is equal to 2 and -2

OpenStudy (anonymous):

the square root of 4 does not equal -2

OpenStudy (anonymous):



OpenStudy (anonymous):

(-2)^2= 4 -2 x -2 = 4

OpenStudy (anonymous):

The domain restricts our range. Because we can't plug in anything lower than 4 into the equation the range is therefor restricted by 4 into infinity. That is why we find out what value we get for y when x=4.